OpenStudy (anonymous):

Write the standard form of the line with the slope of 3 and y-intercept -2. Please help

OpenStudy (godlovesme):

can u write that in slope-intercept first? :3

OpenStudy (godlovesme):

\(\huge\color{violet}{WELCOME~TO~OS~:)}\) @vynguyen1997

OpenStudy (anonymous):

I dont know how to do this problem to write it into a slope intercept. I'm really bad at math.

OpenStudy (godlovesme):

its alright :) 

OpenStudy (anonymous):

so they already give me the y-intercept so i can use -2 for b?

OpenStudy (godlovesme):

yep :) and 3 for m

OpenStudy (anonymous):

y= 3x + -2 is that right?

OpenStudy (godlovesme):

close but + -2 will be -2 so the equation will be y=3x-2 =)

OpenStudy (anonymous):

oh thank you. so my next step is to solve it?

OpenStudy (godlovesme):

yw :) now the next step is to move the 3x to the other side

OpenStudy (anonymous):

so i am guessing it will be y-3x=-2

OpenStudy (anonymous):

is that right?

OpenStudy (godlovesme):

yep good job :D i gave u a medal awesome job!!

OpenStudy (anonymous):

thank you so that is my final answer?

OpenStudy (godlovesme):

yep :D

OpenStudy (anonymous):

omg your so awesome thank you very much. Can you help me with 4 more problems please. sorry for asking because this is my 6th time doing the test :(

OpenStudy (godlovesme):

its all good :) and ty ^.^

OpenStudy (anonymous):

Write the slope-intercept form of the line that passes through the point (1, 0) and is parallel to x - y = 7.

OpenStudy (godlovesme):

can u solve for y in x-y=7 ?

OpenStudy (anonymous):

x-y =7 minus x from both sides then it will come out y= 7-x Is that right?

OpenStudy (anonymous):

is that right?

OpenStudy (godlovesme):

it was supposed to be -y= 7-x

OpenStudy (anonymous):

may i ask why y need to be -y?

OpenStudy (godlovesme):

hold on..

OpenStudy (anonymous):

and is that my final answer for the problem?

OpenStudy (anonymous):

sorry

OpenStudy (godlovesme):